OpenStudy (anonymous):

Combine the like terms and simplify 13p+7-14p-8

OpenStudy (anonymous):

\(13-14=?\)

OpenStudy (anonymous):

13p-14p +7-8

OpenStudy (anonymous):

u see that's where the 13-14 comes from.

OpenStudy (anonymous):

u minus the coefficients.

OpenStudy (anonymous):

@Luis_Rivera ..... u are not meant to provide the answer.

OpenStudy (anonymous):

and let her attempt to do it first.

OpenStudy (anonymous):

13p+7-14p-8 According to the distributive property, for any numbers a, b, and c, a(b+c)=ab+ac and (b+c)a=ba+ca. Here, p is a factor of both 13p and -14p. (13-14)p+7-8 To add integers with different signs, subtract their absolute values and give the result the same sign as the integer with the greater absolute value. In this example, subtract the absolute values of 13 and -14 and give the result the same sign as the integer with the greater absolute value. (-1)p+7-8 Remove the parentheses. -p+7-8 Subtract 8 from 7 to get -1. -p-1